OpenStudy (anonymous):

how to graph x - y = -3 and 2x + y = -3

OpenStudy (austinl):

For this problem, we would solve for x as we did before.

OpenStudy (anonymous):

I got x - y = 3

OpenStudy (anonymous):

first solve for y x - y = -3 x - y + y = -3 + y => x = -3 + y x + 3 = -3 + y +3 x + 3 = y or y = x + 3

OpenStudy (anonymous):

okay:)

OpenStudy (austinl):

Then, you have them in slope intercept form, which you can then graph!

OpenStudy (anonymous):

lets find some points x = 0 (easier to solve) y = 3 so you got the point (0,3) the line passes trhought the point (0,3) slop of 1

OpenStudy (anonymous):

thats it the line that goes from 3rd quadrant to 1st quadrant

OpenStudy (anonymous):

thank you